# Customize-It-3D: High-Quality 3D Creation from A Single Image Using Subject-Specific Knowledge Prior

### v1

@@ Abstract
In this paper, we present a novel two-stage approach that fully utilizes the information provided by the reference image to establish a customized knowledge prior for image-to-3D generation. While previous approaches primarily rely on a general diffusion prior, which struggles to yield consistent results with the reference image, we propose a subject-specific and multi-modal diffusion model. This model not only aids NeRF optimization by considering the shading mode for improved geometry but also enhances texture from the coarse results to achieve superior refinement. Both aspects contribute to faithfully aligning the 3D content with the subject. Extensive experiments showcase the superiority of our method, Customize-It-3D, outperforming previous works by a substantial margin. It produces faithful 360-degree reconstructions with impressive visual quality, making it well-suited for various applications, including text-to-3D creation. Our project page is: https://nnanhuang.github.io/projects/customize-it-3d/
Figure 1: Customize-It-3D has the capability to generate high-fidelity 3D content from just a single image. We present the normal map and novel-view renderings for general objects, highlighting the consistency texture and geometry with remarkable quality at novel views.
1 Introduction
@@
3D creation from a single image remains a profoundly intricate challenge, due to its inherently ill-posed nature, arising from the significant information gap between a single 2D image and the comprehensive 3D spatial and textural properties of the depicted object. Despite its great challenge, recent works [59, 46, 38, 68] have achieved notable results in image-to-3D creation. They draw inspiration from text-to-3D methods [45, 64] which utilize score distillation sampling (SDS) loss to optimize a neural radiance field (NeRF). This SDS loss serves as a guiding principle for novel view synthesis, emulating the way humans utilize priors to infer 3D shape and texture from a single image. While these methods tend to produce promising results, they still exhibit noticeable inconsistencies when compared with the reference view, either in context of the geometry (i.e., multi-face issue) or from the texture perspective (novel views often lack fidelity, presenting smooth and disruptive perticulars).
@@
When crafting intricate 3D models of a specific object, a professional artist indeed need not be equipped with an exhaustive understanding of all the objects in the natural world. Rather, the essential requirement lies in possessing specialized, profound knowledge pertaining to that specific subject at hand. This targeted expertise shall be more instrumental in the meticulous creation of 3D assets. Despite incorporating several adjustments such as CLIP loss and textual inversion considered in [59, 46], the supervision of novel views still relies on a general text-to-image model [49, 51, 53, 3, 18], which has challenge in producing consistent results based solely on text, given the inherent difficulty in crafting textual descriptions to encompass every detail. Consequently, these approaches struggle to maintain consistency and fall short in reconstructing high-fidelity 3D objects. This realization has inspired us to explore strategies for cultivating subject-specific knowledge prior, aiming to achieve high-quality image-to-3D creation.
We propose Customize-It-3D, a novel approach that fully unleashes the geometry/appearance information provided by the reference image to promote customized knowledge prior for 3D generation. Customize-It-3D, like the previous methods mentioned above, adopts a two-stage framework in a coarse-to-fine manner. However, it is guided with subject-specific knowledge prior.
@@
In the first stage, we optimize a neural radiance field to learn an implicit volume representation. The reference view is directly supervised through rendering loss, specifically pixel differences. For novel views, unlike conventional approaches employing general T2I models for SDS loss, we use a subject-specific T2I model. There has been remarkable success in subject-driven image generation [52, 29]. They typically demand at least 3 to 5 images capturing the same subject with varying contexts in order to model the key visual features of the subject. However, our scenario presents a challenge as we possess only a solitary image, which may potentially misguide the model to a trivial solution (mode collapse). To surmount this limitation, we propose to exploit multiple modalities extracted from the reference image, namely depth map, mask image, and normal map. Furthermore, by adjusting the model to be cognizant of the subject’s geometry information, we can supervise NeRF generation in a more fine-grained manner, particularly by taking the shading mode into consideration.
@@
While NeRF is renowned for effectively learning complex geometries, their proficiency in generating intricate texture details is hindered by significant memory demands. Therefore, our objective in the second stage is to enhance visual realism by transforming the coarse NeRF model into point clouds. Conversion to point clouds allows us to have the desirable groundtruth textures on the reference view through projection. Yet we observe that constructing point clouds from depth images as in [59] tends to introduce geometric inaccuracies since NeRF does not store any 3D geometry explicitly (only the density field). In lieu of this, our novel strategy entails an initial step of mesh generation, followed by the established mesh regularization techniques to refine the shape, before sampling points on the mesh surface. In addition, we harness the potential of the aforementioned subject-specific Text-to-Image (T2I) diffusion model, effectively elevating the texture realism, before extending the RGB rendering from the coarse NeRF to point clouds. This augmentation is shown to yield superior texture quality in the final outcomes.
Besides employing the proposed subject-specific 2D diffusion prior, we also utilize a 3D prior to further guide the training process as in [46], compensating the 3D information. Note that this 3D prior takes the reference image as input and thus both the adopted 2D and 3D prior are subject-specific knowledge priors to guide the novel views. To evaluate our method, we collect a benchmark consisting of 100 images. We compare Customize-It-3D with prior state-of-the-arts on RealFusion15 [38], and our Customize100 dataset. Extensive experiments demonstrate our method achieves significant improvement in visual quality in context of accurate geometry and realistic details. In summary, our contributions are:
@@
  • We propose Customize-It-3D, an innovative framework for the generation of high-quality 3D content from a single image by utilizing a subject-specific diffusion prior, which is designed to enhance the personalization of 3D content creation for specific subjects.
@@
  • We introduce a multi-modal DreamBooth model that promotes comprehensive knowledge priors derived from multi-modal images of the subject, prioritizing the faithfulness of the diffusion model to the reference image. This knowledge prior not only facilitates NeRF optimization based on the shading mode but also enhances the texture of the initial coarse results for better refinement.
@@
  • The resulting framework demonstrates state-of-the-art performance in 3D reconstruction using a diverse range of in-the-wild images and images from existing datasets. Furthermore, this approach opens up intriguing possibilities, such as text-to-3D content creation.
2 Related work
@@
Muti-view 3D reconstruction. Early works [1, 55, 14, 54] traditionally require a dense set of input images to accurately deduce geometry by establishing correspondence within the overlapping regions.. The advent of NeRF [39, 36] and its variants [41, 5, 6] have significantly propelled the quality of synthesized novel views through the reconstruction of 3D scenes via implicit neural representations. Subsequent efforts [20, 25, 11, 70] have been dedicated to the facilitation of NeRF optimization from a sparser set of input views. Typically they entail the extraction of per-view features from each input image and aggregate multi-view features for each point along the camera ray. This aggregated information is then decoded to determine density (or Signed Distance Function, SDF) and colors. Diffusion models, which have demonstrated significant advancements in general image synthesis, has emerged as a recent focal point in 3D generation research. Some studies endeavor to directly train 3D diffusion models based on diverse 3D representations, including point clouds [37, 42, 72, 75], meshes [15, 35], neural fields [2, 7, 9, 13, 17, 22, 24, 26, 28, 16, 43, 63, 73]. Others, such as [4, 65, 58, 57] conceptualize 3D generation as an image-to-image translation task and directly generate coherent multiview color images, as exemplified by Zero123 [34]. It’s noteworthy that both approaches require access to multi-view images for training. However, a recurring challenge in this domain lies in the constrained availability of expansive 3D assets datasets. Consequently, most studies have been confined to evaluation on limited categories of shapes or encounter difficulties in extending their models effectively on general objects.
3D generation using knowledge prior. A surge of approaches have been study leveraging large models pretrained on billions of images as knowledge prior to guide the 3D synthesis, circumventing the need for large-scale 3D datasets. One popular task is text-to-3D synthesis, given the exceptional breakthroughs in text-to-image generation. Pioneering works [40, 56] such as DreamFields [21], DreamFusion [45] and SJC [62] have demonstrated success of incorporating NeRF-like models with frozen generative systems for optimization-based 3D synthesis. They fed the rendered 2D images to diffusion models or CLIP model for calculating losses to guide the 3D shape optimization. It sparked numerous follow-up works improve such per-shape optimization scheme in the context of 3D representations [32, 60, 8, 30], sampling schedules [19], and loss design [64], subject-driven editing [48]. An alternative avenue of research [38, 59, 46, 33] seeks to generate 3D digital content from a single image, offering a more precise and controllable approach than text-based methods. The prevailing majority of existing endeavors in this area remain rooted in optimization-based techniques, achieving this task by imposing pixel-wise reconstruction losses on the reference view, in addition to inheriting the SDS loss from text-to-3D methodologies. In comparison, we also follow the DreamFusion pipeline [45], but with a focus on cultivating subject-specific knowledge prior. While Dreambooth3D [48] also leverages a subject-specific diffusion prior, it requires multiple subject images and is designed for text-to-3D editing. In contrast, our approach tackles a more challenging scenario with only one image and introduces a multi-modal diffusion prior, specifically aiming at image-to-3D generation.
@@
Figure 2: We propose a two-stage framework for high-quality 3D creation from a reference image with subject-specific diffusion prior (Sec. 3.1). At the coarse stage, we optimize a NeRF for reconstructing the geometry of the reference image in a shading-aware manner(Sec. 3.2). We further build point clouds with enhanced texture from the coarse stage, and jointly optimize the texture of invisible points and a learnable deferred renderer to generate realistic and view-consistent textures (Sec. 3.3).
3 Method
Generating a personalized 3D object from a single image is a highly challenging task, because it involves disclosing the identity, geometric shape, and texture from limited and incomplete visual information. Therefore, we propose a two-stage coarse-to-fine framework, Customize-It-3D, leveraging carefully cultivated subject-specific knowledge prior to effectively constrain the coherency of 3D object with respect to a particular identity. The proposed framework is illustrated in Figure 2. We will first detail our subject-specific knowledge prior and then introduce our method.
3.1 Subject-Specific Knowledge Prior
@@
Subject-Specific 2D Prior: Multi-modal DreamBooth. Current text-to-image diffusion models possess rich semantic knowledge and 2D image knowledge. DreamFusion [45] is a pioneering work that utilizes the latent knowledge of these diffusion models to guide a 3D representation optimization. However, due to the abundant imagination of 2D diffusion models, it is challenging to precisely control over the generated geometric shapes, textures, and identities using solely text, resulting in discrepancies and thereby causing confusion in guiding 3D reconstructions. This phenomenon tends to produce smoothed textures and imprecise geometries as a result of averaging the inconsistent outcomes. To address this issue, we undertake a fine-tuning process on a pre-trained text-to-image diffusion model using a single reference image. To mitigate the risks of overfitting and mode collapse, we choose to leverage multi-modal images in our approach.
@@
Additionally, we utilize a prior preservation loss, which enables the supervision of model training during fine-tuning with self-generated images $\mathbf{x}_{ori}$ from the original diffusion model. Then our loss function to finetune a diffusion model $\bm{\epsilon}_{\phi}$ is :
$\begin{split}\mathbb{E}_{x,c,\epsilon,\epsilon^{\prime},t}[w_{t}||\bm{\epsilon}_{\phi}{(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,c)-\mathbf{x}||_{2}^{2}}+\\ \lambda w_{t^{\prime}}||\bm{\epsilon}_{\phi}(\alpha_{t^{\prime}}\mathbf{x}_{ori}+\sigma_{t^{\prime}}\epsilon^{\prime},c_{ori})-\mathbf{x}_{ori}||_{2}^{2}],\end{split}$ (1)
where $\mathbf{x}$ is the reference image and $\alpha_{t},\sigma_{t},w_{t}$ control the noise schedule, $c_{ori}$ is the prompt to generate $x_{ori}$. The second term is prior-preservation loss with weight $\lambda$.
In terms of prompt processing, we need an identifier with a weak prior in both the language model and the diffusion model. Therefore, we follow previous work [61] and use the “sks” identifier for stable diffusion [51]. Differently, since we are dealing with a set of images with different modalities, it is imperative not to employ an identical prompt for all images, as this practice can introduce confusion to the model and lead it to erroneously emphasize common distinctive features. Our primary objective is not to model the shared aspects across all modalities but rather to equip the model with the capability to discern and accommodate the distinctions arising from different modalities. Thus, we incorporate additional instructions, and the prompt $c$ for fine-tuning the diffusion model takes the form of “a depth map / normal map / foreground mask / rgb photo of sks [class name]”. The choice of the specific instruction is made in accordance with the given modality during training. In this way, the model is encouraged to perceive the subject from different perspectives, enriching its subject-specific prior.
Subject-Specific 3D Prior: Viewpoint and Image-Conditioned Diffusion Model. Presently, large-scale pre-trained diffusion models are primarily trained on 2D images, lacking substantial 3D-specific knowledge. Notably, the recent Zero-1-to-3 model [34] stands out as it utilizes fine-tuning on a synthetic 3D dataset [10] to develop a viewpoint-conditioned diffusion model. This model takes a reference image along with related external camera parameters as inputs, enabling the generation of novel views of the same subject based on the reference image. Therefore, it serves as our preferred strategy for establishing a subject-specific 3D prior, contributing to improving 3D consistency.
3.2 Coarse Stage: Image to 3D Reconstruction
@@
Subject-specific knowledge prior for novel views. We optimize the neural radiance field such that its multi-view renderings look like high-quality samples using the subject-specific diffusion model. Specifically, let the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v})$ at the viewpoint $\mathbf{v}$, where $\mathcal{G}$ is the differentiable rendering function for NeRF optimization parameterized by $\theta$. We employ the multi-modal DreamBooth model as a 2D prior and follow [45]  [32] to use the score distillation (SDS) loss, which assigns a “score” to the rendered image, guiding the optimization of the 3D model’s parameters $\theta$ towards the direction of higher density regions.
Figure 3: The proposed shading mode-aware guidance largely enhances the 3D geometry in the coarse stage.
Differently, we go beyond this and propose modifying the text prompt based on the specific NeRF shading mode. This modification is aimed at fully harnessing the capabilities of our multi-modal diffusion model to offer precise guidance tailored to different NeRF shading modes, which is facilitated by the inherent multi-modal awareness of our finetuned diffusion model. To elaborate, given the text prompt $y$ that is generated from an image captioning model [31] from the reference image, we adapt the text prompt to “sks normal map of $y$” when the NeRF rendering employs “normal” shading mode and “sks rgb photo of $y$” when the NeRF rendering utilizes “albedo” shading mode. By enforcing consistency between the rendered image aligned with the modified text prompt, the multi-modal diffusion model is able to provide more accurate guidance and thus enhance the quality of the 3D reconstruction, as illustrated in Figure 3. Formally, we use 2D SDS loss as:
@@
where $\mathbf{I}_{m}=\mathcal{G}_{\theta}(\mathbf{v},m)$ and $m$ denotes the shading mode and can take on values of either “normal” or “albedo”. We use the personalized text prompt $y_{m}$ to encode NeRF’s novel view rendering $\mathbf{I}_{m}$ to the noisy latent $z_{t}$ by adding a random Gaussian noise $\bm{\epsilon}$ of a timestep $t$.
In addition, we have the subject-specific 3D prior for novel view guidance, which can be formulated as follows:
$\mathcal{L}_{SS3D}=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\phi}(\mathbf{z_{t};x},t,R,T)-\epsilon)\frac{\partial\mathbf{I}}{\partial\theta}\right],$ (3)
where $R,T$ represent the camera’s rotation and translation, and $\mathbf{x}$ stands for the input reference image. Here, we do not utilize text as guidance but instead use image-based guidance, specifically the relative camera transformations between viewpoints.
The overall supervision is:
@@
Here $\mathrm{Cov(.)}$ denotes covariance and $\mathrm{Var(.)}$ measures standard deviation. We employ this function as our depth regularization to encourage the NeRF output $d$ under the reference view to be close to the depth prior.
Figure 4: Illustrating different point cloud building methods from (1) depth images as in [59] and (2) our method.
@@
Overall training. In summary, the loss in the coarse stage consists of the following parts: $\mathcal{L}_{ref}$, $\mathcal{L}_{SS-SDS}$ and $\mathcal{L}_{depth}$. During training, we follow [59], adopting a progressive training strategy. We start by training only in a partial region of the object’s front view, and then gradually expand the training scope.
3.3 Refine Stage: Neural Texture Enhancement
In the coarse stage, due to the computationally intensive nature of NeRF optimization, we only perform optimization at low resolutions. Moreover, due to the inherent limitations of NeRF, such as the tendency to produce high-frequency artifacts [45], we can only obtain a low-resolution, low-texture-quality 3D model. As a result, we are motivated to explore alternative representations for NeRF, particularly focusing on transforming it into an explicit point cloud. Point clouds have the advantage of allowing direct projection of images, preserving high-quality frontal texture details, and offering opportunities for personalization and enhancement of the projected images, fulfilling the needs of high-quality customized 3D models. In summary, our goal in the refine stage is to retain the geometric structure of the coarse stage, generate a dense point cloud, and optimize the textures not visible in the reference view $\mathbf{v}_{ref}$, as we directly project the frontal reference image $\mathbf{x}$ onto the point cloud. We elaborate several key designs to facilitate the transformation from the coarse to the refine stage.
Point cloud building. Past work has focused on constructing a point cloud from multi-view RGBD images under NeRF rendering [59]. However, due to noise in RGBD images and inaccurate depth estimation, the generated point cloud is very noisy and inaccurate, significantly impairing the 3D geometry in refine stage. To address this, we export the NeRF in coarse stage as a mesh, utilizing mesh regularizations and perform Poisson sampling on the mesh to obtain a dense point cloud $P^{m}$. Figure 4 shows this difference. We opt to a point cloud because it allows personalized diffusion model-based texture enhancement for each viewpoint, resulting in a higher-quality point cloud that better aligns with our customized 3D object generation needs.
Figure 5: The texture details from coarse rendering are largely enhanced using the multi-modal DreamBooth.
Texture projection. However, since NeRF-rendered images from different viewpoints may have overlapping regions, a 3D point may be assigned different colors under different viewpoints [66], leading to texture conflicts. Therefore, we propose an iterative strategy for constructing a clean texture point cloud $P$ from mesh sampled point cloud $P^{m}$. Initially, we trust all points $P^{m}_{ref}$ under reference view $\mathbf{v}_{ref}$ and use a reference mask $M$ to perform color mapping on $P^{m}_{ref}$, obtaining the front-facing perspective of the new point cloud $P_{ref}$. For novel view projection, we aim to avoid introducing points with color conflicts that overlap with $P_{ref}$. Hence, we proceed to project the recently obtained point cloud $P_{ref}$ onto each novel view $\mathbf{v}_{i}$ to generate a new mask denoted as $M^{ref\_proj}_{\mathbf{v}_{i}}$. We then project textures to the points under this novel view only when these points belong to the set difference between the novel view mask $M_{\mathbf{v}_{i}}$ and the reprojected mask $M^{ref\_proj}_{\mathbf{v}_{i}}$. Finally we have a clean point cloud $P=\{p_{ref},p_{1},...,p_{n}\}$ which ensures color mapping is done without introducing conflicts.
Figure 6: Illustrating the textured point cloud construction. We enhance the texture and mask of novel views and iteratively project texture to the point cloud.
Texture and mask enhancement. The texture quality obtained in the coarse stage is suboptimal. Therefore, prior to projecting the reference image $\mathbf{x}$ from the reference view and rendered RGB images $\mathcal{X}=\{\mathbf{x}_{\mathbf{v}_{1}},\mathbf{x}_{\mathbf{v}_{2}},...,\mathbf{x}_{\mathbf{v}_{n}}\}$ under novel views onto the point cloud geometry, we perform a texture enhancement process on all rendered images within $\mathcal{X}$. Specifically, we add noise to the rendered images and leverage the previously fine-tuned multi-modal personalized diffusion model to denoise into a set of pseudo images $\mathcal{X}^{pseudo}$ that possess higher texture quality as well as multi-view consistency. Figure 5 shows that the texture details can be significantly improved by utilizing our multi-modal DreamBooth model.
Simultaneously, due to the issues in the coarse stage of NeRF, the generated masks =$\mathcal{M}=\{M_{\mathbf{v}_{1}},M_{\mathbf{v}_{2}},...,M_{\mathbf{v}_{n}}\}$ by NeRF are not satisfactory. Fine details (for example long and thin parts) in the object structure may be compromised, leading to defects in point cloud texture and geometry initialization due to mask inadequacies. Therefore, we adopt Segment Anything Model (SAM) [27] to generate more refined and accurate masks $\mathcal{M}^{pseudo}$. However, due to the insufficient resolution of RGB images and the noise introduced by SAM, we assess the quality of the generated mask by relying on SAM’s score. Only when the quality surpasses a certain threshold do we consider using this mask; otherwise, we adhere to utilizing masks generated by NeRF. We illustrate our textured point cloud building in Figure 6.

### v3

# High-Quality 3D Creation from A Single Image Using Subject-Specific Knowledge Prior

@@ Abstract
In this paper, we address the critical bottleneck in robotics caused by the scarcity of diverse 3D data by presenting a novel two-stage approach for generating high-quality 3D models from a single image. This method is motivated by the need to efficiently expand 3D asset creation, particularly for robotics datasets, where the variety of object types is currently limited compared to general image datasets. Unlike previous methods that primarily rely on general diffusion priors, which often struggle to align with the reference image, our approach leverages subject-specific prior knowledge. By incorporating subject-specific priors in both geometry and texture, we ensure precise alignment between the generated 3D content and the reference object. Specifically, we introduce a shading mode-aware prior into the NeRF optimization process, enhancing the geometry and refining texture in the coarse outputs to achieve superior quality. Extensive experiments demonstrate that our method significantly outperforms prior approaches.
I INTRODUCTION
The development of robotic models is increasingly constrained by a critical bottleneck: the limited diversity of 3D data available for training. While general image datasets, such as ImageNet[10], contain thousands of object categories, robotics datasets are far less diverse. For instance, ShapeNet[5] and GraspNet[14], two commonly used datasets, include only 55 and 88 object categories. This limitation restricts the ability of robots to effectively interact with a wide range of real-world objects. To address this, there is an urgent need to expand the variety of 3D assets in robotics datasets, enabling more robust and versatile policy models.
At the same time, we also aim to leverage image-to-3D techniques to improve robot interactions with real-world objects. For example, when a robot captures an image of an object using its 2D camera, we can convert that image into a 3D model, thereby enhancing the robot’s ability to interact with and manipulate the object. This approach has the potential to significantly improve robotic performance in complex, dynamic environments.
@@ I INTRODUCTION
However, this task remains deeply intricate due to its inherently ill-posed nature, as there is a substantial domain gap between a single 2D image and the full 3D spatial and textural characteristics of the object. Despite its great challenge, recent works [64, 49, 40, 72] have achieved notable results in image-to-3D creation. They draw inspiration from text-to-3D methods [48, 68] which utilize score distillation sampling (SDS) loss to optimize a neural radiance field (NeRF). While producing promising results, they still exhibit noticeable inconsistencies when compared with the reference view, either in context of the geometry (i.e., multi-face issue) or from the texture perspective (novel views often lack fidelity, presenting smooth and disruptive particulars).
@@
Despite incorporating several adjustments such as CLIP loss and textual inversion considered in [64, 49], the supervision of novel views still relies on a general text-to-image model [52, 54, 56, 3, 19], which has challenge in producing consistent results based solely on text, given the inherent difficulty in crafting textual descriptions to cover every detail. Therefore, these approaches struggle to maintain consistency and fall short in reconstructing high-fidelity 3D objects. This realization has inspired us to explore strategies for cultivating subject-specific knowledge prior, aiming to achieve high-quality image-to-3D.
We employ a two-stage framework, which is a widely recognized way for achieving high-quality image-to-3D. We introduce a subject-specific diffusion prior, customized by the reference image, to fully utilize the geometric and appearance information embedded within the reference image. Building on this prior, we propose shading mode-aware guidance in the NeRF optimization process, and texture enhancement in the refinement stage. We name our approach Customize-It-3D, reflecting the utilization of the proposed customized prior.
@@
In the first stage, we optimize a neural radiance field to learn an implicit volume representation. The reference view is directly supervised through rendering loss, specifically pixel differences. For novel views, unlike conventional approaches employing general T2I models for SDS loss, we use a subject-specific T2I model. There has been remarkable success in subject-driven image generation [55, 30]. They typically demand at least 3 to 5 images capturing the same subject with varying contexts in order to model the key visual features of the subject. However, our scenario presents a challenge as we possess only a solitary image, which may misguide the model to a trivial solution (mode collapse). To surmount this limitation, we propose to exploit multiple modalities extracted from the reference image, namely depth map, mask image, and normal map. More importantly, by adjusting the model to incorporate the subject’s geometric information, we can more precisely guide the generation of NeRF, especially by considering the shading mode in our supervision.
@@
While NeRF is renowned for effectively learning complex geometries, their proficiency in generating intricate texture details is hindered by significant memory demands. Therefore, our objective in the second stage is to enhance visual realism by transforming the coarse NeRF model into point clouds. Conversion to point clouds allows us to have the desirable groundtruth textures on the reference view through projection. Yet we observe that constructing point clouds from depth images as in [64] tends to introduce geometric inaccuracies since NeRF does not store any 3D geometry explicitly (only the density field). In light of this, our strategy entails an initial step of mesh generation, followed by the established mesh regularization techniques to refine the shape, before sampling points on the mesh surface. In addition, we harness the potential of the aforementioned subject-specific diffusion model, to effectively elevate the texture realism, before extending the RGB rendering from the coarse NeRF to point clouds. This augmentation is shown to yield superior texture quality in the final outcomes.
To evaluate our method, we collect a benchmark including 100 images, and compare with baselines on RealFusion [40], and our dataset. In summary, our contributions are:
@@
  • We propose Customize-It-3D, an innovative framework designed to address the critical bottleneck in robotics caused by the limited diversity of 3D data. By generating high-quality 3D content from a single image using a subject-specific diffusion prior, this framework enhances the personalization of 3D asset creation, which is essential for expanding object variety in robotics datasets.
@@
  • We introduce a multi-modal DreamBooth model that promotes comprehensive knowledge priors derived from multi-modal images of the subject, prioritizing the faithfulness of the diffusion model to the reference image. This knowledge prior not only facilitates shading mode-aware NeRF optimization but also enhances the texture for better refinement.
@@
  • The resulting framework demonstrates state-of-the-art performance in 3D reconstruction using a diverse range of in-the-wild images and images from existing datasets.
II Related work
@@
Muti-view 3D reconstruction. Early works [1, 58, 15, 57] traditionally require multiple input images to deduce geometry by establishing correspondence within the overlapping regions. The advent of NeRF [41, 37, 43, 6, 7] have significantly propelled the quality via implicit neural representations. Subsequent efforts [22, 26, 11, 75] have been dedicated to facilitate NeRF optimization from a sparser set of input views. Typically they entail the extraction of per-view features from each input image and aggregate multi-view features for each point. This aggregated information is then decoded to determine density (or Signed Distance) and colors. Diffusion models have emerged as a recent focal point in 3D generation research. Some studies endeavor to directly train 3D diffusion models based on diverse 3D representations, including point clouds [39, 44, 76, 79], meshes [16, 36], neural fields [2, 8, 9, 13, 18, 23, 25, 27, 29, 17, 45, 67, 77]. Others, such as [4, 69, 62, 60] conceptualize 3D generation as an image-to-image translation and directly generate coherent multiview images, as exemplified by Zero123 [34].
3D generation using knowledge prior. Pioneering works [42, 59] such as DreamFields [21], DreamFusion [48] and SJC [65] sparked numerous follow-up works which [40, 64, 49, 33, 66, 70, 38, 33, 61] seeks to generate 3D digital content from a single image, offering a more precise and controllable approach. There are two categories along this direction. One is data-driven approach that trains a feed-forward model with extensible 3D supervision. For instance, LRM [20] adopts a large transformer-based encoder-decoder architecture for learning 3D representations of objects from a single image. DMV3D [73] trains on large scale multi-view image datasets of highly diverse object. Another existing endeavors in this area remain rooted in optimization-based techniques, imposing pixel-wise reconstruction losses on the reference view, in addition to the SDS loss. For example, SyncDreamer [35] adopts a synchronized multiview diffusion to model the joint probability distribution of multiview images for multiview consistency. Magic123 [49] uses 2D and 3D priors simultaneously to generate faithful 3D content from any given image. DreamGaussian [63] designs a generative 3D Gaussian Splatting model with companioned mesh extraction and texture refinement in UV space, aiming at accelerating the optimization process.
In comparison, we follow the optimization-based pipeline [48], but focus on cultivating subject-specific knowledge prior. While Dreambooth3D [51] also leverages customized diffusion prior, it requires multiple subject images and is designed for text-to-3D editing. Instead, our approach tackles a more challenging scenario with only one image and introduces a multi-modal diffusion prior.
@@
Figure 1: We propose a two-stage framework for 3D creation from a reference image with subject-specific diffusion prior (Sec. III-A). At the coarse stage, we optimize a NeRF for reconstructing the geometry of the reference image in a shading-aware manner(Sec. III-B). We further build point clouds with enhanced texture from the coarse stage, and jointly optimize the texture of invisible points and a learnable deferred renderer to generate realistic and view-consistent textures (Sec. III-C).
III Method
In this work, we propose to leverage carefully cultivated subject-specific knowledge prior to effectively constrain the coherency of 3D object to a particular identity to address the need for expanding 3D assets in robotics. The proposed framework is illustrated in Figure 1.
III-A Subject-Specific Knowledge Prior
@@
Subject-Specific Prior: Multi-modal DreamBooth. Current text-to-image diffusion models possess rich semantic knowledge and 2D image knowledge. DreamFusion [48] is a pioneering work that utilizes the latent knowledge of these diffusion models to guide a 3D representation optimization. However, due to the abundant imagination of 2D diffusion models, it is challenging to precisely control over the generated geometric shapes, textures, and identities using solely text, resulting in discrepancies and thereby causing confusion in guiding 3D reconstructions. This phenomenon tends to produce smoothed textures and imprecise geometries as a result of averaging the inconsistent outcomes. To address this issue, we undertake a fine-tuning process on a pre-trained text-to-image diffusion model using a single reference image. To mitigate the risks of overfitting and mode collapse, we choose to leverage multi-modal images in our approach.
@@
Additionally, we utilize a prior preservation loss, which enables the supervision of model training during fine-tuning with self-generated images $\mathbf{x}_{o}$ from the original diffusion model. The loss is,
$\begin{split}\mathbb{E}_{x,c,\epsilon,\epsilon^{\prime},t}[w_{t}||\bm{\epsilon}_{\phi}{(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,c)-\mathbf{x}||_{2}^{2}}+\\ \lambda w_{t^{\prime}}||\bm{\epsilon}_{\phi}(\alpha_{t^{\prime}}\mathbf{x}_{o}+\sigma_{t^{\prime}}\epsilon^{\prime},c_{o})-\mathbf{x}_{o}||_{2}^{2}],\end{split}$ (1)
where $\mathbf{x}$ is the reference image and $\alpha_{t},\sigma_{t},w_{t}$ control the noise schedule, $c_{o}$ is the prompt to generate $x_{ori}$ and c is the prompt with a unique identifier for reference image.
In terms of prompt processing, we need an identifier with a weak prior in both the language model and the diffusion model. Therefore, we follow previous work [46] and use the “sks” identifier for stable diffusion [54]. Differently, since we are dealing with a set of images with different modalities, it is imperative not to employ an identical prompt for all images, as this practice can introduce confusion to the model and lead it to erroneously emphasize common distinctive features. Our primary objective is not to model the shared aspects across all modalities but rather to equip the model with the capability to discern and accommodate the distinctions arising from different modalities. Thus, we incorporate additional instructions, and the prompt $c$ for fine-tuning the diffusion model takes the form of “a depth map / normal map / foreground mask / rgb photo of sks [class name]”. The choice of the specific instruction is made in accordance with the given modality during training. In this way, the model is encouraged to perceive the subject from different perspectives.
III-B Coarse Stage: Image to 3D Reconstruction
@@
Figure 2: An example of a Lego castle. Point cloud building results from (1) depth images [64] and (2) our method.
Subject-specific knowledge prior for novel views. We optimize the neural radiance field using the subject-specific diffusion model. Specifically, let the rendered image $\mathbf{I}=\mathcal{G}_{\theta}(\mathbf{v})$ at the viewpoint $\mathbf{v}$, where $\mathcal{G}$ is the differentiable rendering function for NeRF optimization parameterized by $\theta$. We employ the multi-modal DreamBooth model as a 2D prior and follow [48]  [32] to use the score distillation (SDS) loss, which assigns a “score” to the rendered image, guiding the optimization of the 3D model’s parameters $\theta$ towards the direction of higher density regions.
Differently, we go beyond this and propose modifying the text prompt based on the specific NeRF shading mode. This modification is aimed at fully harnessing the capabilities of our multi-modal diffusion model to offer precise guidance tailored to different NeRF shading modes, which is facilitated by the inherent multi-modal awareness of our finetuned diffusion model. To elaborate, given the text prompt $y$ that is generated from an image captioning model [31] from the reference image, we adapt the text prompt to “sks normal map of $y$" when the NeRF rendering employs “normal" shading mode and “sks rgb photo of $y$" when the NeRF rendering utilizes “albedo" shading mode. By enforcing consistency between the rendered image aligned with the modified text prompt, the multi-modal diffusion model is able to provide more accurate guidance and thus enhance the quality of the 3D reconstruction. Formally, we use 2D SDS loss as:
@@
where $\mathbf{I}_{m}=\mathcal{G}_{\theta}(\mathbf{v},m)$ and $m$ denotes the shading mode and can take on values of either “normal" or “albedo". We use the personalized text prompt $y_{m}$ to encode NeRF’s novel view rendering $\mathbf{I}_{m}$ to the noisy latent $z_{t}$ by adding a random Gaussian noise $\bm{\epsilon}$ of a timestep $t$.
In addition, we leverage a 3D prior for novel view guidance, thanks to the recent Zero-1-to-3 model [34]. The 3D loss can be formulated as follows:
$\nabla_{\theta}\mathcal{L}_{SS3D}=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\phi}(\mathbf{z_{t};x},t,R,T)-\epsilon)\frac{\partial\mathbf{I}}{\partial\theta}\right],$ (3)
where $R,T$ represent the camera’s rotation and translation, and $\mathbf{x}$ stands for the input reference image. The overall supervision is:
@@
Here $\mathrm{Cov(.)}$ denotes covariance and $\mathrm{Var(.)}$ measures standard deviation. We employ this function as our depth regularization to make the NeRF output $d$ under the reference view close to the depth prior.
@@
Overall training. In summary, the loss in the coarse stage consists of $\mathcal{L}_{ref}$, $\mathcal{L}_{SS-SDS}$ and $\mathcal{L}_{depth}$. During training, we follow [64], adopting a progressive training strategy. We start by training only in a partial region of the object’s front view, and then gradually expand the training scope.